OpenStudy (anonymous):

going to ask this again, please I need help, this is my last one for tonight. 4x-8/y+2 divided by y-2/y^2-4

OpenStudy (gw2011):

I assume this is the problem: [(4x-8)/(y+2)]/[(y-2)/(y^2-4)] [(4x-8)/(y+2)][(y^2-4)/(y-2)] [(4x-8)/(y+2)][(y-2)(y+2)/(y-2)] The answer is: 4x-8 = 4(x-2)

OpenStudy (anonymous):

\[\left(\begin{matrix}4x-8 \\ y+2\end{matrix}\right)\div \left(\begin{matrix}y-2 \\ y ^{2}-4\end{matrix}\right)\]

OpenStudy (gw2011):

O.K. The answer I gave is for the problem you just stated.

OpenStudy (anonymous):

ok thank you :)

OpenStudy (gw2011):

You're welcome